A field microphone used at a soccer game has a parabolic cross section and is 24 inches wide at the opening. The focus is 16 inches from the vertex. Find the depth of the microphone.
Answer by lwsshak3(11628)

A field microphone used at a soccer game has a parabolic cross section and is 24 inches wide at the opening. The focus is 16 inches from the vertex. Find the depth of the microphone.
***
Consider the microphone like a parabola that opens right with vertex at the origin.
Its basic equation: y^2=4px
For given parabolic microphone:
p=16
4p=64
equation: y^2=64x
12^2=64x
144=64x
x=144.64/64
x=2.25
depth of the microphone=2.25 inches